# Targeting Alignment: Extracting Safety Classifiers of Aligned LLMs

### v2

@@ 5.3 Safety Classifier
Start point of the classifier. We focus on the structures $f_{1,1+\delta}$, i.e., contiguous sets of decoders starting with the first layer. This allowed to locate the end of the safety classifier by evaluating the performance of the corresponding candidates. However, it may be that the classifier does not start at the first layer. Therefore, studying structures $f_{i,i+\delta}$ with $i>1$ may be of interest to refine the extraction but introduce a new dimension of complexity. Since the input space of decoder $i$ must match the output space of decoder i1i-1, new techniques need to be introduced to allow the creation of candidates that truly use the information of the structure. For example, a candidate classifier based on $f_{i,i+\delta}$ should not use information from $f_{1,i-1}$ or learn new information outside the mapping from structure representations to classification. We leave such explorations to future work.
@@
Labels. Part of our approach involves obtaining the classifications from the LLM. This involves two steps: generating the output from the LLM, then mapping this output text to a class, using the classification rule $\mathcal{R}$ defined in Section 3.3. LLMs are deployed in a way that allows nondeterministic outputs, leading to probabilities of refusal and compliance. To ensure reproducibility and reduce computational costs, we set the temperature to 0. However, studying the temperature and how it influences the probabilities of labels would lead to a more precise understanding of safety classifier and its decision boundary. We leave this study to future work.
@@
Finding optimal depth. From an adversarial perspective, finding the position of the safety classifier is crucial, as it can be expensive to run the attack on various candidates. A first strategy to pinpoint the most likely optimal depth would be to follow the procedure of Section 3.1, i.e., using the silhouette score. However, the adversary may be constrained in resources and have limited GPU memory. In that case, multiple strategies are possible, such as starting at the maximum candidate size for the GPU memory and lowering it or doing a binary search on the candidate size.
6 Related Work
6.1 Jailbreak and Representations
@@ 6.2 Adversarial Settings
From an attack perspective, there are three ways to use the component: ablation [60], intervention [21], or optimization of the input to attack the component [26]. Ablation corresponds to disabling the component during generation, preventing the model from classifying the safety of the input. Interventions operate by adding a perturbation to the embedding at the position of the component. If the component is a classifier, the perturbation corresponds to the direction orthogonal to its decision boundary.
In this work, we extract a surrogate classifier which can be attacked directly with a heuristic-free adversarial objective. While interventions can be applied to the candidates, they may not lead to feasible input perturbations (i.e., it may not be possible to find a sequence of tokens that represent the intervention), which limits the characterization of alignment failures in adversarial settings.
6.3 Pruning
In addition to studying their safety, an active area of research on LLMs is pruning, i.e., reducing the size of the model while maintaining most of its capabilities [13]. For LLMs, pruning can be divided into three axes: depth, width, and length pruning. Depth pruning focuses on removing layers of the model, width pruning aims to reduce the size of the projection matrices, and length pruning focuses on cutting tokens of the input stream and avoiding redundant computations. As those axes are orthogonal, they can be combined to obtain better results [43]. The rise of security concerns led to the intersection of pruning and security. Recent work has shown that safety alignment is inherently low rank and can be isolated [48] and that pruning can offer higher safety [15]. Although these techniques could achieve a similar objective, they generally involve modifying weights from the selected parts of the model. As we aim to extract the safety classifier, minimizing the information lost on its decision boundary by having as little training as possible is crucial, which is why we chose the least invasive method (linear probing). This is also the most efficient way to extract the classifier, as only the necessary weights are trained, i.e., those that map the representation to the label.
7 Conclusion
In this paper, we hypothesized that alignment embeds a safety classifier within LLMs, and we introduced techniques to extract best candidate classifiers for it. We empirically verified the agreement with the models of the best candidates in both benign and adversarial settings. We also demonstrated how the extracted classifier can be used to systematically study alignment security, as the candidate classifiers substantially improved both the efficiency (lower runtime and VRAM usage) and the efficacy (higher success rate) for attackers and defenders. Finally, while we studied alignment in a safety context, other failures of alignment such as hallucination, dishonesty, bias, and lack of fairness could also benefit from our approach.
@@ A.1 Confusion Matrices
Figure 13: ASR on the candidates and transferability rate to the LLM after applying the white-box attack on the candidates for safe samples.
A.2 Attacking Safe Inputs
As with any classification problem, mitigating false negatives (jailbreaks) comes at the cost of higher false positives (over-refusal). This effect was seen in the lower classification performance on OR-Bench (designed to test this phenomenon) in Section 4.2. In this section, we focus on attacking safe inputs to create a refusal.
Setup. We also augment the dataset by adding targets corresponding to a refusal. Different aligned models have different refusal answers depending on how they were trained. However, we saw limited impact on the choice of the target string on the performance of the attack and kept the same target for all prompts and models: “I cannot fulfill your request. I’m just an AI”.
Safe ASR. Similarly to Section 4.3, the attack is applied to the models, this time with safe inputs (with the objective of making the model refuse). Table IV reports the ASR for all settings. We first see that the highest ASR corresponds to Llama 2 (99%), which has the lowest ASR on unsafe inputs (12% to 22%). In contrast, attacking Qwen 2.5 produced a high ASR for unsafe and safe inputs. Finally, Gemma 1 has shown medium ASR on unsafe inputs, also exhibited for safe inputs.
@@ A.2 Attacking Safe Inputs
Results. 13(a) and 13(b) report the ASR on candidate classifiers and the transferability rate to the LLM, for safe inputs. We notice the same trends on the transferability rate with a peak for most models with an exception for Llama 2. Furthermore, for OR-Bench, the ASR on the candidate classifiers corresponds exactly to the transferability rate, implying that the success of the attack is entirely determined by how well it performs on the candidate classifier. This finding aligns with the performance in adversarial setting in 8(b): Llama 2 candidate classifiers have the most stable performance on adversarial examples from the LLM compared to other models.
A.3 Other Models
@@
Models. We studied three other models: Llama-3.1-8B-Instruct [31], Mistral-7B-Instruct-v0.3 [19], and Zephyr_RMU [22].
@@
Results. Table V reports their accuracy and F1 score on the two datasets. All models achieve an F1 score lower than 0.8 on OR-Bench, translating to a weaker performance according to the setup. Figure 14 and Figure 15 show the performance of the candidates in benign settings, with the same setup as Section 4.2. We see that the trend is similar in 14(a) with a convergence after a certain candidate size. However, the F1 score is lower than the other models studied. Further, 14(b) shows a different trend, with a decrease at the middle of the model. Figure 15 shows the cross-dataset experiment for the other models studied.
@@
Possible causes. The Mistral model is described as a ”quick demonstration” that ”does not have any moderation mechanisms” [^5] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 in the ”Limitations” section [19]. Llama 3, similarly to Llama 2, did go through alignment. However, the tone of the output was adjusted, specifically for refusal outputs. Therefore, detecting the classification made by the model is more challenging. Zephyr RMU is a model in which unsafe information has been removed through unlearning [22], therefore the lower performance is due to the way labels are assigned: in our experimental setup, we assign labels based on whether the model complies with the input prompt or not. Thus, since Zephyr RMU underwent unlearning, it may still comply with unsafe input prompts with a safe output. We argue that models that are made more secure through unlearning are orthogonal to the scope of the paper: this method likely removes part of the safety classifier alongside the unsafe knowledge, preventing its extraction.

### v3

@@ V-C Safety Classifier
Start point of the classifier. We focus on the structures $f_{1,1+\delta}$: contiguous sets of decoders starting at the first layer. This allowed to locate the end of the safety classifier by evaluating the performance of the corresponding candidates. However, the classifier may not start at the beginning. Therefore, studying structures $f_{i,i+\delta}$ with $i>1$ may be of interest to refine the extraction but introduce a new dimension of complexity. Since the input space of decoder $i$ must match the output space of decoder i1i-1, new techniques need to be introduced to allow the creation of candidates that truly capture the information of the structure. For example, a candidate classifier based on $f_{i,i+\delta}$ should not use information from $f_{1,i-1}$. We leave such explorations to future work.
@@
Labels. Part of our approach involves obtaining the classifications from the LLM. This involves two steps: generating the output from the LLM, then mapping this output text to a class, using the classification rule $\mathcal{R}$ defined in Section III-C. LLMs are deployed in a way that allows nondeterministic outputs, leading to probabilities of refusal and compliance. To ensure reproducibility and reduce computational costs, we set the temperature to 0 in our experiments. Studying the temperature and its influence over the probabilities of labels would lead to a more precise understanding of safety classifier and its decision boundary. We leave this study to future work.
@@
Finding optimal depth. From an adversarial perspective, finding the position of the safety classifier is crucial, as it can be expensive to run the attack on various candidates. A first strategy to pinpoint the most likely optimal depth would be to follow the procedure of Section III-A, i.e., using the silhouette score. However, the adversary may be constrained in resources and have limited GPU memory. In that case, multiple strategies are possible, such as starting at the maximum candidate size for the GPU memory and lowering it or doing a binary search on the candidate size.
VI Related Work
VI-A Jailbreak and Representations
@@ VI-B Adversarial Settings
From an attack perspective, there are three ways to use the component: ablation [65], intervention [22], or optimization of the input to attack the component [27]. Ablation disables the component during generation, preventing the model from classifying the safety of the input. Interventions operate by adding a perturbation to the embedding at the position of the component. In the case of a classifier, the perturbation corresponds to the direction orthogonal to its decision boundary.
In this work, we extract a surrogate classifier which can be attacked directly with a heuristic-free adversarial objective. While interventions can be applied to the candidates, they may not lead to feasible input perturbations, limiting the characterization of alignment failures.
VI-C Pruning
In addition to studying their safety, an active area of research on LLMs is pruning, i.e., reducing the size of the model while maintaining most of its capabilities [14]. For LLMs, pruning can be divided into three axes: depth, width, and length pruning. Depth pruning focuses on removing layers of the model, width pruning aims to reduce the size of the projection matrices, and length pruning focuses on cutting tokens of the input stream and avoiding redundant computations. As those axes are orthogonal, they can be combined to obtain better results [46]. The rise of security concerns led to the intersection of pruning and security. Recent work has shown that safety alignment is inherently low rank and can be isolated [52] and that pruning can offer higher safety [16]. Although these techniques could achieve a similar objective, they generally involve modifying selected parts of the model. As we aim to extract the safety classifier, minimizing the information lost on its decision boundary by having as little training as possible is crucial, which is why we chose the least invasive and most efficient method (linear probing).
VII Conclusion
In this paper, we hypothesized that alignment embeds a safety classifier within LLMs, and we introduced techniques to extract best candidate classifiers for it. We empirically verified the agreement with the models of the best candidates in both benign and adversarial settings. We also demonstrated how the extracted classifier can be used to systematically study alignment security, as the candidate classifiers substantially improved both the efficiency (lower runtime and VRAM usage) and the efficacy (higher success rate) for attackers and defenders. Finally, while we studied alignment in a safety context, other failures of alignment such as hallucinations or lack of fairness could also benefit from our approach.
@@ Acknowledgments
The authors thank Chaowei Xiao and Rahul Chatterjee as well as all reviewers for their helpful comments on previous iterations of the work.
@@ A-A Confusion Matrices
Figure 13: ASR on the candidates and transferability rate to the LLM after applying the attack on the candidates for safe samples.
A-B Attacking Safe Inputs
As with any classification problem, mitigating false negatives (jailbreaks) comes at the cost of higher false positives (over-refusal). This effect was seen in the lower classification performance on OR-Bench (designed to test this phenomenon) in Section IV-B. In this section, we focus on attacking safe inputs to create a refusal.
Setup. We also augment the dataset by adding refusal targets. Different aligned models have different refusal answers depending on how they were trained. However, we saw limited impact on the choice of the target string on the performance of the attack and used the same target for all prompts and models: “I cannot fulfill your request. I’m just an AI”.
Safe ASR. Similarly to Section IV-C, the attack is applied to the models, this time with safe inputs (with the objective of making the model refuse). Table IV reports the ASR for all settings. We first see that the highest ASR corresponds to Llama 2 (99%), which has the lowest ASR on unsafe inputs (12% to 22%). In contrast, attacking Qwen 2.5 produced a high ASR for unsafe and safe inputs. Finally, Gemma 1 shows medium ASR on unsafe inputs, also exhibited for safe inputs.
@@ A-B Attacking Safe Inputs
Results. 13(a) and 13(b) report the ASR on candidate classifiers and the transferability rate to the LLM, for safe inputs. We notice the same trends on the transferability rate with a peak for most models with an exception for Llama 2. Furthermore, for OR-Bench, the ASR on the candidate classifiers corresponds exactly to the transferability rate, implying that the success of the attack is determined by how well it performs on the candidate classifier. This finding aligns with the performance in adversarial settings in 8(b): Llama 2 candidate classifiers have the most stable performance on adversarial examples from the LLM compared to other models.
A-C Other Models
@@
Models. We studied three other large language models: Llama-3.1-8B-Instruct [33], Mistral-7B-Instruct-v0.3 [20], and Zephyr_RMU [23].
@@
Results. Table V reports their accuracy and F1 score on the two datasets. All models achieve an F1 score lower than 0.8 on OR-Bench, translating to a weaker performance according to the setup. Figure 14 and Figure 15 show the performance of the candidates in benign settings, with the same setup as Section IV-B. We see that the trend is similar in 14(a) with a convergence after a certain candidate size. However, the F1 score is lower than the other models studied. Further, 14(b) shows a different trend, with a decrease at the middle of the model. Figure 15 shows the cross-dataset experiment for the other models studied.
@@
Possible causes. The Mistral model is described as a “quick demonstration” that “does not have any moderation mechanisms” [^5] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 in “Limitations”. [20]. Llama 3, similarly to Llama 2, did go through alignment. However, the tone of the output was adjusted, specifically for refusal outputs. Therefore, detecting the classification made by the model is more challenging. Zephyr RMU is a model in which unsafe information has been removed through unlearning [23], therefore the lower performance is due to the way labels are assigned: in our experimental setup, we assign labels based on whether the model complies with the input prompt or not. Thus, since Zephyr RMU underwent unlearning, it may still comply with unsafe input prompts with a safe output. We argue that models that are made more secure through unlearning are orthogonal to the scope of the paper: this method likely removes part of the safety classifier alongside the unsafe knowledge, preventing its extraction.